An iron rod is 2.58m long at 0°C. Calculate the length of a brass rod at 0°C, if the difference between the lengths of the two rods must remain the same at all temperature. Linear expansivity of iron = 1.2 × 10^-5k^-1. Linear expansivity of brass = 1.9 × 10^-5k^-1
1
Expert's answer
2019-03-05T09:43:08-0500

By the definition of the linear thermal expansion we have:


"\Delta L_{iron} = \alpha_{iron} L_{0, iron} \Delta T_{iron},""\Delta L_{brass} = \alpha_{brass} L_{0, brass} \Delta T_{brass},"

here, "\Delta L_{iron}", "\Delta L_{brass}" are the difference in the lengths of the copper and aluminum rods after the change in the temperature, respectively; "L_{0, iron}", "L_{0, brass}" are the length of the copper and aluminum rods before the change in the temperature (at 0°C), respectively; "\alpha_{iron}", "\alpha_{brass}" are the coefficients of linear expansion for the copper and aluminum rods, respectively; "\Delta T_{iron}", "\Delta T_{brass}" are the change in temperature, respectively.

Since the difference between the lengths of the two rods must remain the same at all temperature, we can write:


"\Delta T_{iron} = \Delta T_{brass} = \Delta T,""\Delta L_{iron} = \Delta L_{brass}."

Then, we get:


"\alpha_{iron} L_{0, iron} \Delta T = \alpha_{brass} L_{0, brass} \Delta T,""L_{0, brass} = \dfrac{\alpha_{iron} L_{0, iron}}{\alpha_{brass}} = \dfrac{1.2 \cdot 10^{-5} K^{-1} \cdot 2.58 m}{1.9 \cdot 10^{-5} K^{-1}} = 1.63 m."

Answer:

"L_{0, brass} = 1.63 m."